# Fine-Grained Cross-View Localization via Local Feature Matching and Monocular Depth Priors

### v1

@@ 3.3 Model Supervision
We follow Xia and Alahi (2025) to provide supervision on both the pose and correspondences. For camera pose, we adopt the Virtual Correspondence Error (VCE) loss from Xia and Alahi (2025). Specifically, we define a set of $N_{v}$ virtual points $\mathbf{P}_{v}$ in a 2D metric space, where the hyperparameter $l$ controls the spatial extent of the space. We apply both the ground-truth transformation, $\mathbf{R}_{\text{gt}},\mathbf{t}_{\text{gt}}$, and the estimated transformation, $\mathbf{R},\mathbf{t}$, to these virtual points. The loss $\mathcal{L}_{\text{VCE}}$ then minimizes the mean Euclidean distance between the corresponding transformed points,
@@
When metric depth is available during training, we use the ground-truth pose to find, for the ground points, their corresponding aerial points and compute an infoNCE loss Oord et al. (2018), $\mathcal{L}_{\text{G2S}}$, to encourage these correspondences. We also find, for the aerial points, their corresponding ground points. However, multiple ground points may exist at different heights but share similar BEV locations. We select the ground point closest to the projected location and compute the $\mathcal{L}_{\text{S2G}}$, using only ground points outside a defined local neighborhood as negatives. Therefore, we do not limit the matching between aerial and ground images to be one-to-one. The detailed formulation is provided in the Appendix. The total loss is then a weighted combination of the VCE loss and infoNCE losses, $\mathcal{L}=\mathcal{L}_{\text{VCE}}+\beta(\mathcal{L}_{\text{G2S}}+\mathcal{L}_{\text{S2G}})/2$.
@@ 4.1 Datasets and Evaluation Metrics
KITTI Geiger et al. (2013) provides forward-facing ground-level images with a limited field of view. Aerial images for KITTI were collected by Shi and Li (2022), who also split the dataset into Training, Test1, and Test2 subsets. Test1 includes images from the same region as the Training set, while Test2 contains images from a different region. We refer to Test1 and Test2 as the same-area and cross-area test sets, respectively. KITTI provides ground-truth depth maps, which are commonly used in training depth foundation models Yang et al. (2024b, a). We use the metric depth predictions from Yang et al. (2024b) for our experiments on KITTI.
VIGOR Zhu et al. (2021) is a widely used cross-view localization dataset, containing ground-level panoramas and aerial images from four U.S. cities. It provides two splits: same-area and cross-area. In the same-area split, training and test images are collected from all four cities, while in the cross-area split, the training and test sets come from two different cities. VIGOR does not have ground-truth depth. During training, we adopt a recent metric depth model Piccinelli et al. (2025), which has shown strong performance on outdoor panoramic images. At test time, we report results using both this metric depth and relative depth from various methods Wang and Liu (2024); Jiang et al. (2021); Wang et al. (2022). Additionally, we include results in the Appendix where both training and testing are performed using only relative depth.
Metrics: We report the mean and median localization and orientation errors. For KITTI, we further decompose localization errors into longitudinal and lateral components based on the driving direction, and report the percentage of samples within the defined error thresholds.
@@ 4.2 Implementation Details
For both datasets, we use AdamW Loshchilov and Hutter (2019) with a learning rate of $1\times 10^{-4}$. Training is conducted on a single H100 GPU with a batch size of 80 for VIGOR and 224 for KITTI. We follow the setting from Xia and Alahi (2025): the temperature parameter $\tau$ in matching score computation is set to $0.1$, the aerial feature map $\mathbf{F}_{A}$ is resized to generate $41\times 41$ aerial points, $\mathcal{L}_{\text{VCE}}$ uses $N_{v}=10\times 10$ points, with $l$ set to 5 m, and $N=1024$ correspondences are sampled for pose estimation. For $\mathcal{L}_{\text{S2G}}$, negatives are 1 m away in planar distance from the projected location. We use $\beta=1$ for VIGOR and $\beta=0.1$ for KITTI. During testing on VIGOR, we use RANSAC Fischler and Bolles (1981) for outlier filtering. RANSAC is not used on KITTI, as we did not observe any improvement in localization accuracy. When using metric depth, we apply a maximum depth threshold of 35 m for VIGOR and 40 m for KITTI, setting the matching score in $\hat{M}^{*}$ to zero for points exceeding this threshold. For relative depth predictors, we apply a fixed initial scale (identified by visually inspecting a few examples) to all relative depth maps and then use the threshold to filter out matches corresponding to sky and distant objects.
@@ 4.3 Quantitative Results
Table 1: KITTI test results. Best in bold. During both training and testing, our method utilizes metric depth predictions from DepthAnythingv2 Yang et al. (2024b).
Methods $\downarrow$ Loc. (m) $\uparrow$ Lateral ($\%$) $\uparrow$ Long. ($\%$) $\downarrow$ Orien. () $\uparrow$ Orien. ($\%$)
@@
Same-area GGCVT Shi et al. (2023) - - 76.44 98.89 23.54 62.18 - - 99.10 100.00
CCVPE Xia et al. (2023) 1.22 0.62 97.35 99.71 77.13 97.16 0.67 0.54 77.39 99.95
HC-Net Wang et al. (2024b) 0.80 0.50 99.01 99.73 92.20 99.25 0.45 0.33 91.35 99.84
DenseFlow Song et al. (2024) 1.48 0.47 95.47 99.79 87.89 94.78 0.49 0.30 89.40 99.31
@@
Ours 1.11 0.76 99.60 100.00 65.86 98.04 1.97 1.43 36.68 92.84
Cross-area GGCVT Shi et al. (2023) - - 57.72 91.16 14.15 45.00 - - 98.98 100.00
CCVPE Xia et al. (2023) 9.16 3.33 44.06 90.23 23.08 64.31 1.55 0.84 57.72 96.19
HC-Net Wang et al. (2024b) 8.47 4.57 75.00 97.76 58.93 76.46 3.22 1.63 33.58 83.78
DenseFlow Song et al. (2024) 7.97 3.52 54.19 91.74 23.10 61.75 2.17 1.21 43.44 89.31
[…]
Ours 5.51 2.97 90.97 99.22 13.70 65.06 3.32 2.12 26.03 80.68
KITTI: As shown in Tab. 1, our method significantly outperforms previous state-of-the-art methods on the cross-area test set, while achieving slightly higher localization error on the same-area test set. When test samples are from a different area than training, our method lowers the mean localization error by 26% (from 7.45 m to 5.51 m) compared to previous state-of-the-art Xia and Alahi (2025), demonstrating stronger generalization capability. Our orientation error is slightly higher than that of the baselines. Notably, models that directly regress orientation Shi et al. (2023); Xia et al. (2023); Song et al. (2024); Wang et al. (2024b) perform well on KITTI, as most forward-facing ground images align with the road direction, an implicit prior that these models can easily learn. In contrast, our method computes orientation analytically from local feature correspondences, which makes it harder to exploit this prior but offers greater interpretability.
VIGOR: Overall, our method demonstrates strong and consistent performance across all settings, including known/unknown orientations and same-/cross-area scenarios, see Tab. 2. For example, while FG2 achieves lower localization error when the orientation is known, our method significantly outperforms it in the more challenging unknown orientation setting. This highlights our model’s practical applicability in real-world scenarios where camera orientation is not available. Compared to the previous state-of-the-art under unknown orientation Xia et al. (2023), our method achieves comparable localization accuracy: slightly higher mean error on the same-area test set, but lower mean error on the cross-area test set. Notably, our method consistently achieves the lowest mean orientation error across all settings, demonstrating its effectiveness in jointly estimating both location and orientation.
[…]
Ori. Methods Same-area Cross-area
@@
Known SliceMatch Lentsch et al. (2023) 5.18 2.58 - - 5.53 2.55 - -
CCVPE Xia et al. (2023) 3.60 1.36 - - 4.97 1.68 - -
GGCVT Shi et al. (2023) 4.12 1.34 - - 5.16 1.40 - -
DenseFlow Song et al. (2024) 3.03 0.97 - - 5.01 2.42 - -
HC-Net Wang et al. (2024b) 2.65 1.17 - - 3.35 1.59 - -
@@
Ours-BiFuse++ 3.17 1.69 - - 3.56 2.02 - -
Ours-UniFuse 3.14 1.67 - - 3.54 2.01 - -
Unknown SliceMatch Lentsch et al. (2023) 6.49 3.13 25.46 4.71 7.22 3.31 25.97 4.51
CCVPE Xia et al. (2023) 3.74 1.42 12.83 6.62 5.41 1.89 27.78 13.58
DenseFlow Song et al. (2024) 4.97 1.90 11.20 1.59 7.67 3.67 17.63 2.94
[…]
Ours 3.94 1.78 9.54 2.00 4.23 2.09 11.67 2.21
[…]
Ours-BiFuse++ 4.12 1.94 10.13 2.31 4.43 2.26 12.27 2.47
Ours-UniFuse 4.03 1.89 9.81 2.21 4.36 2.21 12.10 2.43
[…]
Second, we evaluate a more practical scenario where only a relative depth model Wang et al. (2022); Jiang et al. (2021); Wang and Liu (2024) is available in inference. We use different relative depth models in a plug-and-play manner with our estimated correspondences, without retraining or finetuning our model for relative depth. As shown in the bottom two rows of the known and unknown orientation entries in Tab. 2, using relative depth leads to only a modest increase, less than 0.2 m, in localization error, regardless of whether the evaluation is conducted in same-area or cross-area scenarios. This flexibility in incorporating different depth predictors highlights the practicality and robustness of our method for real-world deployment.
@@
4.4 Qualitative Results
Local feature matching: The core of our accurate and robust localization lies in the reliability of our correspondence matching between ground and aerial images. Among existing methods, only our approach and FG2 Xia and Alahi (2025) are capable of producing such local feature correspondences, providing superior interpretability. However, FG2 fails when the orientation of the camera is unknown.
We show in Fig. 3 examples of our local feature matching. As shown in (a) and (b), our method accurately matches road markings, such as yellow grid lines and text painted on the road surface, across ground and aerial views. In (c) and (d), our method uses streetlights as landmarks to anchor positions and establish correspondences between views. Interestingly, in (e), our method matches an overhead road sign in the ground view to its shadow in the aerial image, demonstrating that the matching is not solely appearance-based but reflects strong semantic understanding. Finally, (f) highlights that, without constraining ground keypoints to lie on object tops, our method establishes reasonable matches between building facades in the ground view and their rooftops in the aerial view.
Scale recovery: Next, we demonstrate that when a relative depth predictor Wang et al. (2022) is used in inference, our method can jointly estimate the camera pose and the scale between the relative depth and the aerial metric space. Fig. 4 shows the alignment by overlaying the rotated, translated, and scaled BEV points onto the aerial image using the estimations. The transformed points show consistent alignment with the aerial imagery, highlighting the accuracy of both the estimated pose and scale.
Figure 4: BEV points overlaid on the aerial image after applying the predicted rotation, translation, and scale transformations.
@@ 4.5 Ablation Study
As our method does not introduce new deep learning modules and most hyperparameters are adopted from prior state-of-the-art Xia and Alahi (2025), our ablation study focuses on our key design choices. In particular, we study: (1) the coordinate assignment strategy (Sec.3.2), where we compare keeping all points within the defined depth range versus retaining only the topmost points along height; and (2) the impact of ignoring scale in Procrustes alignment (which brings our formulation to the orthogonal Procrustes used in Xia and Alahi (2025); Barroso-Laguna et al. (2024)). As shown in Tab. 3, explicitly selecting the topmost points does not improve localization accuracy. The side view of an object’s top does not necessarily exhibit the most consistent semantics with the aerial image. Therefore, enforcing the use of topmost points may hinder cross-view localization. Regarding the scale-awareness, when the scale is not considered during training and testing, the model must learn feature matching that aligns strictly with the metric depth predictions. Since metric depth is imperfect, allowing the model to also estimate scale based on the matched correspondences leads to more accurate and robust local feature matching. Moreover, without modeling the scale, the method cannot support relative depth predictors Wang and Liu (2024); Wang et al. (2022); Jiang et al. (2021), since the orthogonal Procrustes assumes that all coordinates lie in the same scale space.
Table 3: Ablation study on VIGOR same-area validation set with unknown orientation. Best in bold.

### v2

@@ 3.3 Model Supervision
We follow Xia & Alahi (2025) to provide supervision on both the pose and correspondences. For camera pose, we adopt the Virtual Correspondence Error (VCE) loss. Specifically, we define a set of $N_{v}$ virtual points $\mathbf{P}_{v}$ in a 2D metric space, where the hyperparameter $l$ controls the spatial extent of the space. We apply both the ground-truth transformation, $\mathbf{R}_{\text{gt}},\mathbf{t}_{\text{gt}}$, and the estimated transformation, $\mathbf{R},\mathbf{t}$, to these virtual points. The loss $\mathcal{L}_{\text{VCE}}$ then minimizes the mean Euclidean distance between the corresponding transformed points,
@@
When metric depth is available during training, we use the ground-truth pose to find, for the ground points, their corresponding aerial points and compute an infoNCE loss (Oord et al., 2018), $\mathcal{L}_{\text{G2S}}$, to encourage these correspondences. We also find, for the aerial points, their corresponding ground points. However, multiple ground points may exist at different heights but share similar BEV locations. We select the ground point closest to the projected location and compute the $\mathcal{L}_{\text{S2G}}$, using only ground points outside a defined local neighborhood as negatives. The detailed formulation is provided in the Appendix. The total loss is then a weighted combination of the VCE loss and infoNCE losses, $\mathcal{L}=\mathcal{L}_{\text{VCE}}+\beta(\mathcal{L}_{\text{G2S}}+\mathcal{L}_{\text{S2G}})/2$.
@@ 4.1 Datasets and Evaluation Metrics
KITTI (Geiger et al., 2013) provides forward-facing ground-level images with a limited field of view. Aerial images for KITTI were collected by Shi & Li (2022), who also split the dataset into Training, Test1, and Test2 subsets. Test1 includes images from the same region as the Training set, while Test2 contains images from a different unseen region. We refer to Test1 and Test2 as the same-area and cross-area test sets, respectively. We use the metric depth predictions from DepthAnythingV2 (Yang et al., 2024b) for our experiments on KITTI.
VIGOR (Zhu et al., 2021) is a widely used cross-view localization dataset, containing ground-level panoramas and aerial images from four U.S. cities. It also provides same-area and cross-area splits. In the same-area split, training and test images are collected from all four cities, while in the cross-area split, the training and test sets come from two different cities. During training, we adopt a recent metric depth model, Unik3D (Piccinelli et al., 2025), which has shown strong performance on outdoor panoramic images. At test time, we report results using both Unik3D and relative depth models (Wang & Liu, 2024; Jiang et al., 2021; Wang et al., 2022). Additionally, we include results in the Appendix where both training and testing are performed using only relative depth.
Metrics: We report the mean and median localization and orientation errors. Following Xia et al. (2023), we train and test with both known and unknown orientations on VIGOR, and apply orientation noise sampled from $\pm 10^{\circ}$ or $\pm 180^{\circ}$ on KITTI. For KITTI, we also decompose localization errors into longitudinal and lateral components based on the driving direction and report the percentage of samples within the defined error thresholds.
@@ 4.2 Implementation Details
On both datasets, we use AdamW (Loshchilov & Hutter, 2019) with a learning rate of $1\times 10^{-4}$. Training is conducted on a single H100 GPU with a batch size of 80 for VIGOR and 224 for KITTI. We follow the setting from Xia & Alahi (2025): the temperature parameter $\tau$ in matching score computation is set to $0.1$, the aerial feature map $\mathbf{F}_{A}$ is resized to generate $41\times 41$ aerial points, $\mathcal{L}_{\text{VCE}}$ uses $N_{v}=10\times 10$ points, with $l$ set to 5 m, and $N=1024$ correspondences are sampled for pose estimation. For $\mathcal{L}_{\text{S2G}}$, negatives are 1 m away in planar distance from the projected location. We use $\beta=1$ for VIGOR and $\beta=0.1$ for KITTI. RANSAC (Fischler & Bolles, 1981) is applied on VIGOR, but omitted on KITTI as it did not yield improvements in localization accuracy. When using metric depth, we apply a maximum depth threshold of 35 m for VIGOR and 40 m for KITTI, setting the matching score of points exceeding this threshold to zero. For relative depth predictors, we apply a fixed initial scale (identified by visually inspecting a few examples) to all relative depth maps and then use the threshold to filter out matches corresponding to sky and distant objects.
@@ 4.3 Quantitative Results
Table 1: KITTI test results. Best in bold. The ‘ori.’ column shows orientation noise used in training and testing, uniformly sampled within $\pm 10^{\circ}$ or $\pm 180^{\circ}$.
Ori. Methods $\downarrow$ Loc. (m) $\uparrow$ Lateral ($\%$) $\uparrow$ Long. ($\%$) $\downarrow$ Orien. () $\uparrow$ Orien. ($\%$)
@@
Cross-area $\pm 10^{\circ}$ GGCVT - - 57.72 91.16 14.15 45.00 - - 98.98 100.00
CCVPE 9.16 3.33 44.06 90.23 23.08 64.31 1.55 0.84 57.72 96.19
HC-Net 8.47 4.57 75.00 97.76 58.93 76.46 3.22 1.63 33.58 83.78
DenseFlow 7.97 3.52 54.19 91.74 23.10 61.75 2.17 1.21 43.44 89.31
$\text{FG}^{2}$ 7.31 4.15 37.89 85.65 21.98 60.77 3.62 2.37 23.03 77.84
Ours 5.60 3.01 45.29 92.43 27.01 68.26 3.32 2.12 26.03 80.68
$\pm 180^{\circ}$ SliceMatch 14.85 11.85 24.00 72.89 7.17 33.12 23.64 7.96 31.69 31.69
CCVPE 13.94 10.98 23.42 60.46 11.81 42.12 77.84 63.84 3.14 14.56
Ours 11.71 9.11 13.64 50.88 14.04 50.73 55.18 33.23 2.53 12.78
Same-area $\pm 10^{\circ}$ GGCVT - - 76.44 98.89 23.54 62.18 - - 99.10 100.00
CCVPE 1.22 0.62 97.35 99.71 77.13 97.16 0.67 0.54 77.39 99.95
HC-Net 0.80 0.50 99.01 99.73 92.20 99.25 0.45 0.33 91.35 99.84
DenseFlow 1.48 0.47 95.47 99.79 87.89 94.78 0.49 0.30 89.40 99.31
$\text{FG}^{2}$ 0.75 0.51 95.81 99.66 92.50 99.05 0.93 0.66 67.27 98.91
Ours 1.13 0.77 93.59 99.97 71.51 98.17 1.97 1.43 36.68 92.84
$\pm 180^{\circ}$ SliceMatch 7.96 4.39 49.09 98.52 15.19 57.35 4.12 3.65 13.41 64.17
CCVPE 6.88 3.47 53.30 85.13 25.84 68.49 15.01 6.12 8.96 42.75
Ours 1.85 1.31 62.18 97.80 59.37 97.48 9.70 6.17 9.46 41.64
KITTI: We compare our method with both global descriptor-based methods, CCVPE (Xia et al., 2023) and SliceMatch (Lentsch et al., 2023), and geometry transformation-based methods, GGCVT (Shi et al., 2023), HC-Net (Wang et al., 2024b), DenseFlow (Song et al., 2024), and FG2 (Xia & Alahi, 2025). As shown in Tab. 1, our method sets a new state-of-the-art in mean and median localization errors on the cross-area test under both $\pm 10^{\circ}$ and $\pm 180^{\circ}$ orientation noise. On the same-area test, although our method has higher error under $\pm 10^{\circ}$ noise, it shows a large improvement over the previous state-of-the-art method in the more challenging $\pm 180^{\circ}$ setting, reducing the mean error from 6.88 m to 1.85 m. For orientation prediction, our method is less accurate than the state of the art. Since our method computes orientation analytically from local feature correspondences, it cannot take full advantage of the prior that most ground images in KITTI are aligned with the road direction. Under $\pm 10^{\circ}$ noise, our performance is comparable to FG2, while under $\pm 180^{\circ}$ noise, our method outperforms CCVPE but falls short of SliceMatch.
VIGOR: Overall, our method demonstrates strong and consistent performance in both localization and orientation estimation on the same-area and cross-area test sets (see Tab. 2). For instance, while FG2 attains lower localization error when the orientation is known, our method significantly outperforms it in the more challenging unknown orientation setting. Compared to the previous state of the art under unknown orientation (Xia et al., 2023), our method achieves comparable localization accuracy, showing slightly higher mean error on the same-area test set but lower mean error on the cross-area test set. Notably, for panoramic images, the richer matchable information between ground and aerial views leads to substantial gains, with our method achieving the lowest mean orientation error in both same-area and cross-area tests.
@@
Ori. Methods Cross-area Same-area
@@
Unknown SliceMatch 7.22 3.31 25.97 4.51 6.49 3.13 25.46 4.71
CCVPE 5.41 1.89 27.78 13.58 3.74 1.42 12.83 6.62
DenseFlow 7.67 3.67 17.63 2.94 4.97 1.90 11.20 1.59
$\text{FG}^{2}$ 10.02 8.14 31.41 5.45 8.95 7.32 15.02 2.94
Ours 4.23 2.09 11.67 2.21 3.94 1.78 9.54 2.00
Ours-BiFuse++ 4.43 2.26 12.27 2.47 4.12 1.94 10.13 2.31
Ours-UniFuse 4.36 2.21 12.10 2.43 4.03 1.89 9.81 2.21
Known SliceMatch 5.53 2.55 - - 5.18 2.58 - -
CCVPE 4.97 1.68 - - 3.60 1.36 - -
GGCVT 5.16 1.40 - - 4.12 1.34 - -
DenseFlow 5.01 2.42 - - 3.03 0.97 - -
HC-Net 3.35 1.59 - - 2.65 1.17 - -
$\text{FG}^{2}$ 2.41 1.37 - - 1.95 1.08 - -
Ours 3.43 1.90 - - 3.06 1.59 - -
@@
Second, we evaluate a practical scenario where only a relative depth model (Wang et al., 2022; Jiang et al., 2021; Wang & Liu, 2024) is available at inference. We plug in different relative depth models with our estimated correspondences, without retraining or finetuning. As shown in the bottom two rows of the known and unknown orientation entries in Tab. 2, relative depth increases localization error by less than 0.2 m. This flexibility in depth predictors highlights the practicality and robustness of our method for real-world deployment, especially when a compact depth predictor is required.
@@
4.4 Interpretability
Local feature matching: Since our pose is computed analytically from the estimated correspondences, these correspondences directly reflect localization quality, an interpretability advantage absent in most prior cross-view localization methods. Although FG2 also estimates correspondences, its localization fails when the orientation is unknown, whereas our method remains accurate.
As shown in Fig. 3 (a) and (b), when road markings are repetitive, e.g., on highways, our method leverages other landmarks such as overhead road signs and streetlights for localization, demonstrating that the matching is not solely appearance-based but also reflects strong semantic understanding. In (c), the partially occluded ‘bus’ marking is difficult even for humans to identify, yet our model matches it correctly. In (d), both streetlights and road markings serve as anchors for localization, while (e) shows that without restricting matches to object tops, our method aligns the upper facades of buildings in the ground view with rooftops in the aerial view. Finally, (f) shows a failure case, though interestingly our method predicts a more reasonable position aligned with the visible path, whereas the ground truth appears erroneous. More VIGOR and KITTI results are in the Appendix.
Outlier detection: The ability to estimate correspondences makes our method well-suited for outlier detection using RANSAC. For each test sample, we record the number of inlier correspondences and use it for outlier detection. As shown in Fig. 4, the pose error decreases sharply as the inlier ratio rises from 10% to 50%, followed by a slower decline after 50%. This trend reveals a strong negative correlation between inlier ratio and pose error, consistent with the expectation that more accurate correspondences yield more accurate pose estimates.
Figure 4: Outlier detection using RANSAC on VIGOR same/cross-area test sets.
Ground-aerial layout alignment: With a relative depth predictor (Wang et al., 2022), our method estimates both the camera pose and the scale between relative depth and the aerial metric space. As shown in Fig. 5, the rotated, translated, and scaled ground layout can be overlaid onto the aerial image, providing a clear and interpretable visual cue of localization quality. For example, (a) and (b) show precise alignment of the projected intersection and building with those in the aerial view, corresponding to accurate localization. Interestingly, (c) also shows good alignment, revealing an error in the ground-truth location: the vehicle should be in front of the zebra crossing rather than on it. (d) shows a localization failure, which is easily identifiable from the misalignment. The strong correlation between alignment and localization quality offers a powerful tool for interpretability.
Refer to caption(a)
Refer to caption(b)
Refer to caption(c)
Refer to caption(d)
Figure 5: Ground layout overlaid on the aerial image after applying the predicted rotation, translation, and scale transformations. The alignment directly reflects localization quality: the first three examples show successful localization, while the last one illustrates a failure case. Notably, the alignment in example (c) helped us to identify the error in ground truth location.
@@ 4.5 Ablation Study
Our main ablation study focuses on two key design choices: (1) the coordinate assignment strategy (Sec. 3.2), comparing using all points within the depth range versus only the topmost points along height; and (2) the effect of ignoring scale in Procrustes alignment (reducing our formulation to the orthogonal Procrustes of Xia & Alahi (2025); Barroso-Laguna et al. (2024)). As shown in Tab. 3, explicitly selecting the topmost points does not improve localization accuracy. The side view of an object’s top does not necessarily exhibit the most consistent semantics with the aerial image. Therefore, enforcing the use of topmost points may hinder cross-view localization. For scale-awareness, excluding scale forces the model to align features strictly to metric depth predictions, which are imperfect. Allowing the model to also estimate scale from correspondences yields more accurate and robust matching. Moreover, without modeling scale, the method cannot support relative depth predictors, as orthogonal Procrustes assumes that all coordinates lie in the same scale space.
Table 3: Ablation study on VIGOR same-area validation set with unknown ori. Best in bold.